OpenStudy (solomonzelman):

I will risk continuing this without shamil98.

OpenStudy (solomonzelman):

I learned so far the rule of power and rule of product doing derivatives. Can someone esle continue tutoring me?

OpenStudy (kc_kennylau):

Have you learnt the chain rule?

OpenStudy (solomonzelman):

No

OpenStudy (kc_kennylau):

The chain rule: \[\frac {dy}{dx}=\frac {dy}{du}\cdot\frac{du}{dx}\]

OpenStudy (kc_kennylau):

For example, y=tan(x^2). Try to find \(\dfrac{dy}{dx}\) using the above formula with \(\large u=x^2\) :)

OpenStudy (solomonzelman):

I never learned this in a dx/dy form, I guess the risk turned out a bad way.

OpenStudy (kc_kennylau):

oh, what is the notation that you learnt?

OpenStudy (kc_kennylau):

I'll try to convert the formula to your notation

OpenStudy (solomonzelman):

\[nx^{x-1}\] like \[x^4~'=4x^3\] AND \[f'g+g'f\]

OpenStudy (solomonzelman):

I guess shamil98 would be my best option....

OpenStudy (kc_kennylau):

\[y'(x)=y'(u)\times u'(x)\] y(x) means y as a function of x.

OpenStudy (solomonzelman):

what does u represent?

OpenStudy (kc_kennylau):

A function of x, for example,

OpenStudy (solomonzelman):

YEah, I am barely able to find regular derivatives, hold on.....

OpenStudy (kc_kennylau):

Let u=cos(x). -ln(cos(x))' = -ln(u)' * u' = -1/u * -sin(x) = -1/cos(x) * -sin(x) = tan(x)

OpenStudy (solomonzelman):

I can';t find the previous q, b.c my prof isn't working.

OpenStudy (solomonzelman):

This is what I was left off. http://openstudy.com/users/shamil98#/updates/52b94eb1e4b01cdca4494f8a

OpenStudy (kc_kennylau):

ok, do you know the derivatives of the trigonometric functions?

OpenStudy (solomonzelman):

I've never done that

OpenStudy (kc_kennylau):

ok, let me teach you two basic derivatives of trigonometric functions: \[\sin(x)'=\cos(x)\]\[\cos(x)'=-\sin(x)\]

OpenStudy (kc_kennylau):

Let's practice what you learnt :) Try to find the derivative of \(\tan(x)\) using the quotient rule :)

OpenStudy (solomonzelman):

A;right, I was about to say nvm.... I need to know the definitions first using easier examples, then do harder and harder. If you go through that question, you will see that I am incapable of deriving tan(x)

OpenStudy (kc_kennylau):

But @shamil98 just taught you the quotient rule...

OpenStudy (solomonzelman):

Got disconnected.... but not how to apply it to trigonometric function though.

OpenStudy (solomonzelman):

Yeah, I'll wait for shamil. kennylau, ty for your help though

OpenStudy (kc_kennylau):

no :/ I can help you :/

OpenStudy (solomonzelman):

Sure?

OpenStudy (kc_kennylau):

ok, find the derivative of \(\dfrac{x^2+1}{x+3}\) using the quotient rule :)

OpenStudy (solomonzelman):

\[2x(x+3)+1(x^2+1)\] right?

OpenStudy (solomonzelman):

\[f'g+g'f\]

OpenStudy (kc_kennylau):

That is the product rule :)

OpenStudy (solomonzelman):

The which one is the quotient?

OpenStudy (kc_kennylau):

\[(f/g)'=(f'g-g'f)/g^2\]

OpenStudy (solomonzelman):

Oh, that part is the unclear part, I stopped like right before that, or at least meant to.

OpenStudy (kc_kennylau):

So let me give you the formulae: \[\begin{array}{rcl}(f+g)'&=&f'+g'\\(f-g)'&=&f'-g'\\(fg)'&=&f'g+g'f\\\left(\dfrac fg\right)'&=&\frac{f'g-g'f}{g^2}\end{array}\]

OpenStudy (kc_kennylau):

as well as: \[\Large(x^n)'=nx^{n-1}\]

OpenStudy (solomonzelman):

Yeah, the last think I do know, it's the power rule. and I get the rest too.

OpenStudy (kc_kennylau):

So do you still wanna do my question? :)

OpenStudy (solomonzelman):

Hold on, I need to get some staff down.

OpenStudy (kc_kennylau):

ok

OpenStudy (kc_kennylau):

Continue to wait for @shamil98 coz I'm in GMT+8 and I need to sleep now :)

OpenStudy (solomonzelman):

Pk, I'll continue later, I need rest.

OpenStudy (kc_kennylau):

What is your timezone? :)

OpenStudy (solomonzelman):

11:13 AM US, eastern coast.